# Representative electricity price profiles for European day-ahead and intraday spot markets

### v1

@@ 5   Conclusions
We present a method to construct electricity price scenarios based on an averaging and scaling procedure using historical price data. Our approach is capable of generating both DA and ID scenarios and the cumulative deviation of a [DA, ID] pair can be adjusted to zero-mean deviation, if comparability is needed. Our method is constructed to be simple and intuitive, to have few adjustable parameters, and thus be easily reproducible, while at the same time providing sufficiently representative price profiles for future case studies on DSM.
We have applied our method to historical DA and ID data from EPEX spot for the year 2023. There, we have generated multiple price scenarios with varying magnitude of price fluctuation. The comparison to historical data confirmed that our proposed method is capable of providing representative price profiles of adjustable variance for both DA and ID markets simultaneously. Moreover, the constructed profiles automatically satisfy a set of basic Requirements 1 to 4. Consequently, the tedious and non-deterministic procedure of selecting a representative historical profiles while ensuring the desired profile properties hold is circumvented.
@@
Clearly, some advanced scenario generation methods also consider higher-order statistical moments in their generation procedure. However, these approaches are inherently more involved and result in a greater number of adjustable parameters. In general, both historical data based as well as complex scenario forecasting methods have been rather inaccurate at predicting the long-term evolution of prices, as this evolution is subject to policy and political changes. Hence, we focus on approaches to reflect the current character of the electricity price on spot markets. Since we are interested in an intuitive and deterministic method rather than covering all possible price profile shapes, we accept to trade such a flexibility for the desired simplicity of the proposed averaging-based method.
@@
Acknowledgements. The authors gratefully acknowledge the financial support of the Kopernikus project SynErgie by the Federal Ministry of Education and Research (BMBF), and the project supervision by the project management organization Projektträger Jülich (PtJ). The authors gratefully acknowledge the financial support of and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)—333849990/GRK2379 (IRTG Hierarchical and Hybrid Approaches in Modern Inverse Problems). Moreover, we thank Sonja Germscheid and Eike Cramer for constructive feedback on the manuscript.
Supplementary information
Chrysanthi Papadimitriou, Jan C Schulze, and Alexander Mitsos
 Abstract. Herein, we provide supplementary material to the listed publication.
Extreme & Moderate DA & ID profiles
(a) Moderate
(b) Extreme
Fig. 1: Moderate ($\beta=1.00$, $\gamma=1.00$) and Extreme ($\beta=1.85$, $\gamma=1.77$) single-day DA and ID price profile for year 2023.
(a) Moderate
(b) Extreme
Fig. 2: Moderate ($\beta=1.00$, $\gamma=1.00$) and Extreme ($\beta=1.76$, $\gamma=2.32$) single-week DA and ID price profile for year 2023.
Best fits of historical single-day and single-week profiles
(a) DA market
(b) ID market
Fig. 3: Extreme single-day price profiles compared to the best fit historical day in 2023.
(a) DA market
(b) ID market
Fig. 4: Extreme single-week price profiles compared to the best fit historical week in 2023.
Fig. 5: Extreme DA single-week price profiles compared to one of the best fitted historical weeks in 2023.
Data of single-day DA and ID profiles
Table 1: Profiles of hourly DA electricity prices.
Hour DA price in €/MWh Hour DA price in €/MWh
$k$ Nominal Moderate Extreme $k$ Nominal Moderate Extreme
1 78.81 84.03 74.61 13 58.92 70.50 49.63
2 72.84 79.97 67.11 14 54.99 67.82 44.69
3 69.47 77.68 62.89 15 63.81 73.83 55.77
4 68.52 77.04 61.69 16 80.34 85.07 76.53
5 74.05 80.80 68.64 17 103.81 101.05 106.02
6 93.27 93.88 92.78 18 133.73 121.42 143.61
7 115.06 108.71 120.15 19 158.45 138.24 174.67
8 120.63 112.5 127.15 20 150.24 132.65 164.35
9 112.18 106.75 116.54 21 129.49 118.53 138.29
10 94.45 94.68 94.27 22 113.27 107.49 117.91
11 80.35 85.08 76.55 23 99.93 98.41 101.15
12 69.27 77.54 62.63 24 88.33 90.51 86.57
Table 2: Profiles of 15 min ID electricity prices.
Quarter ID price in €/MWh Quarter ID price in €/MWh
$q$ Nominal Moderate Extreme $q$ Nominal Moderate Extreme
1 105.11 97.78 98.88 49 83.31 83.24 72.13
2 86.65 88.13 81.84 50 59.63 70.87 50.28
3 69.46 79.15 65.99 51 48.22 64.91 39.75
4 52.32 70.20 50.18 52 28.78 54.75 21.82
5 90.66 89.28 83.56 53 35.71 57.75 26.91
6 73.83 80.49 68.03 54 45.66 62.95 36.08
7 65.65 76.22 60.48 55 53.25 66.92 43.09
8 57.75 72.09 53.19 56 65.72 73.43 54.59
9 75.31 80.73 68.27 57 16.36 49.04 12.00
10 65.99 75.86 59.67 58 52.29 67.81 45.15
11 63.81 74.72 57.66 59 77.41 80.93 68.32
12 62.63 74.10 56.57 60 103.50 94.56 92.39
13 60.35 72.76 54.15 61 20.45 53.79 21.29
14 62.29 73.78 55.94 62 69.7 79.52 66.72
15 69.40 77.49 62.50 63 99.36 95.02 94.09
16 76.46 81.18 69.02 64 129.15 110.58 121.57
17 55.09 70.89 51.14 65 42.19 68.86 49.18
18 64.02 75.55 59.38 66 89.87 93.77 93.17
19 81.40 84.64 75.42 67 125.09 112.17 125.65
20 95.86 92.19 88.75 68 162.54 131.73 160.20
21 58.88 75.91 61.06 69 79.43 93.05 93.52
22 81.30 87.62 81.74 70 115.21 111.74 126.53
23 103.84 99.40 102.53 71 149.51 129.66 158.17
24 121.73 108.75 119.04 72 198.27 155.13 203.15
25 93.37 97.38 100.14 73 143.59 130.48 160.96
26 112.29 107.26 117.60 74 143.09 130.22 160.50
27 130.14 116.58 134.06 75 167.54 142.99 183.05
28 120.20 111.40 124.90 76 202.96 161.50 215.73
29 156.87 131.43 160.58 77 199.54 158.41 209.83
30 118.76 111.52 125.43 78 150.58 132.83 164.66
31 111.31 107.63 118.55 79 139.01 126.79 153.99
32 92.53 97.82 101.23 80 131.18 122.7 146.77
33 149.08 126.02 150.58 81 187.53 148.85 191.83
34 116.79 109.16 120.79 82 136.91 122.41 145.13
35 95.71 98.15 101.35 83 119.83 113.49 129.38
36 69.39 84.39 77.06 84 104.97 105.72 115.66
37 139.50 118.22 135.82 85 158.76 131.26 159.88
38 105.77 100.60 104.71 86 121.80 111.95 125.78
39 80.90 87.60 81.76 87 104.83 103.08 110.12
40 49.30 71.09 52.61 88 78.47 89.31 85.80
41 117.6 104.54 110.91 89 137.52 118.05 135.82
42 87.21 88.67 82.87 90 109.54 103.43 110.01
43 70.87 80.13 67.81 91 93.70 95.16 95.40
44 42.27 65.19 41.42 92 67.40 81.42 71.14
45 100.50 93.86 91.45 93 120.31 107.22 116.08
46 72.70 79.34 65.80 94 99.44 96.32 96.82
47 61.04 73.24 55.03 95 78.06 85.15 77.10
48 38.55 61.5 34.29 96 57.24 74.27 57.89

### v2

@@ 5.1 Scenarios
5.2 Comparison to historical data
For the sake of comparison, we complement the constructed profiles by historical profiles that closely match each of the calculated profiles. Thereby, we show that the constructed profiles do not only artificially represent the averaged historical data but are even similar to real historical price profiles. Such a historic profile may thus be considered close to the center point in the averaging metric. We retrieve the corresponding historical day by minimizing the mean absolute deviation between the historical and a constructed scenario for DA and ID simultaneously. The resulting profiles are visually juxtaposed to the constructed ones. In the single-day case, we solve:
$\min_{i}\Big(\sum_{k=1}^{24}|\overline{DA}_{k}^{\beta}-DA_{i,k}|+\frac{1}{4}\sum_{q=1}^{96}|\overline{ID}_{q}^{\gamma}-ID_{i,q}|\Big)\,.$ (14)
A single-week historical profile is identified analogously. The minimization problem is solved by first calculating the absolute price differences between the constructed daily profiles and each of the historical DA and ID profiles, as given in Eq. (14). Then, the objective value is calculated for each of the historical days, and the minimum value is selected. The corresponding day represents the historical time-series data that best aligns with the constructed DA and ID profiles simultaneously.
Figure 7 contrasts the Nominal single-day profiles and the historical day with the best fit. While the inner hour variance of the Nominal ID profile is more pronounced during early and late hours, all profiles follow similar trends and are overall comparable. Moreover, inserting Eq. (2) into Eq. (14) reveals that the historic profile in Fig. 7 indeed reflects the day which deviates the least from the rest historical days, i.e., the cluster center in this sense. Next, Fig. 8 depicts the Unscaled profiles and the historical best fit. Since here all single-week profiles have the same historical week as the best fit, we only show the plot for the DA Nominal profile in Fig. 9. The rest of the figures for profile matching with historical data are given in the SI.
We raise attention to the higher mismatch of the two profiles in Fig. 9 for the last hours of the week and we note that during the weekdays (from 0 h up to 120 h) the price pattern is well-defined due to a consistent daily cycle in electricity demand (i.e., prices rise when the workday begins, peak in the afternoon, and decline in the evening as demand shifts to residential use) (Knittel and Roberts, 2005). This does not apply to the weekend (from 120 h up to 168 h) when demand is more unpredictable. We conclude that the Unscaled and Nominal profile represent the historical data closely. For the Nominal profile, the fitted historical day merely serves for validation, whereas for the Unscaled profile the historical day corresponds to the most representative day in the historical data set.
Lastly, we discuss a limitation of our approach regarding more extreme scenarios. Different from the close alignment of the Unscaled and Nominal profiles with historical data, solving the fitting problem in Eq. (14) for the Extreme profile did not lead to satisfactory results. This is explained by the extreme deviations, i.e., outliers, observed in the historical data, especially in the ID price. These outliers, despite showing high variance, when composing the Extreme profiles, exhibit different patterns than the averaged ones of the constructed (scaled) profiles. Nevertheless, when fitting historical DA and ID profiles separately rather than jointly, we were able to find similar historical profiles (see figures in SI). We conclude that while a more extreme variance scaling is generally valid, an alternative scenario generation method should be consulted if extreme profiles with abnormal patterns are sought.
6  Case study and method comparison
We apply our scenario generation method to a DSM scheduling case study and compare to scheduling results when employing price scenarios obtained using literature methods. As elaborated in the introductory section, we aim for short-horizon price scenarios that characterize longer periods sufficiently well to enable substantiated generalization and thus decision-making, e.g., process design assessment. Thereby, we target case studies with complex process models, for which long scheduling horizons are computationally prohibitive (Schulze et al., 2024).
We use the DA and ID price scenarios of the suggested Nominal and Unscaled cases and compare to scenarios generated by clustering methods. To this end, we examine the economic generalization error from a generated scenario to the detailed results obtained for one year of real (historical) price data. We then compare the generalization errors from the different price scenarios.
6.1 Problem description
We investigate chlor-alkali electrolysis (CAE), which is a classical DR candidate due to high electricity demand (Roh et al., 2019; Brée et al., 2020). To facilitate scheduling flexibility, we include a downstream product storage tank. Given a fixed daily production volume, we aim to minimize electricity costs over a scheduling horizon. To this end, we consider the three scheduling setups where (i) only the DA spot market is considered, (ii) DA and ID market are included simultaneously, and (iii) a two-stage program considering a DA market in the first stage and ID scheduling subsequently. Setup (ii) represents the idealistic case of an exact ID price forecast on the day ahead. On the other hand, setup (iii) accounts for the chronological order of participation in the two markets.
The scheduling program is formulated as a linear program (LP). In all three cases, we solve deterministic scheduling LPs of the generic form:
$\min_{\bm{u}_{t}\in\Omega}\hskip 4.30554pt\frac{1}{T}\sum_{t=1}^{T}\Phi_{t}(\bm{x}_{t},\bm{u}_{t})$ (15a)
$\mathrm{s.t.}\hskip 17.22217pt$ $\bm{f}(\bm{x}_{t},\bm{x}_{t-1},\bm{u}_{t})=\bm{0}\,,\;\bm{x}_{0}=\hat{\bm{x}}\,,\;t\in\{1,...,T\}\,$, (15b)
$\bm{c}(\bm{x}_{t},\bm{x}_{t-1},\bm{u}_{t})\leq\bm{0}\,,\;t\in\{1,...,T\}\,$, (15c)
$\bm{x}_{k}=\bm{x}_{0}\,,\;k=96,192,...,T\,$, (15d)
wherein $t$ is the time index of quarter-hourly discretization steps on the scheduling horizon, and $T=96$ for a single-day horizon and T=35 040T=35\,040 for a full year. The scheduling degrees of freedom (DOFs), $\bm{u}_{t}\in\Omega$, i.e., the traded energy volume(s) $P_{t}$, are constrained to an admissible set $\Omega\subseteq\mathbb{R}^{n_{u}}$. The states $\bm{x}_{t}\in\mathbb{R}^{n_{x}}$ describe the quater-hourly production volumes $F_{t}$, and mass production rate $M_{t}$, and the energy storage $S_{t}$ (i.e., power integrated over the 15 min between t1t-1 and $t$). Equation (15b), wherein $\bm{f}:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u}}\rightarrow\mathbb{R}^{n_{x}}$, represents the linear process model. The dependence on $\bm{x}_{t-1}$ is due to the integrating storage system. Equation (15c), wherein $\bm{c}:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u}}\rightarrow\mathbb{R}^{n_{c}}$, represents the linear path constraints and ramping constraints on $P_{t}$, and the linear constraints on $M_{t}$, $F_{t}$ and $S_{t}$. The ramping constraints reflect that the process cannot realize arbitrarily large steps in the production rate. Linear constraints on the power purchases from the DA market reassure compliance with the market limitations, i.e., power can be bought up to a maximum allowance and not sold to the DA market. Finally, Equation (15d) represents a periodic constraint on storage at the end of each day, and Equation (15b) gives the fixed initial state $\hat{\bm{x}}\in\mathbb{R}^{n_{x}}$. The detailed equations and parametric values of the process model and problem constraints are taken from Germscheid et al. (2022).
To calculate the running electricity cost $\Phi_{t}$, we follow the notation of electricity prices, $DA_{\lceil t/4\rceil}$ and $ID_{t}$, from the preceding sections. Correspondingly, the purchased DA and ID energy volumes are denoted by $P^{\mathrm{DA}}_{t}$ and $P^{\mathrm{ID}}_{s,t}$. The available scheduling DOFs as well as the calculation of $\Phi_{t}$ differ across the three cases:
  1. (i)
    $\Phi_{t}^{(\mathrm{i})}=DA_{\lceil t/4\rceil}\cdot P^{\mathrm{DA}}_{t}$ and $\bm{u}_{t}=P^{\mathrm{DA}}_{t}$.
  2. (ii)
    $\Phi_{t}^{(\mathrm{ii})}=DA_{\lceil t/4\rceil}\cdot P^{\mathrm{DA}}_{t}+ID_{t}\cdot P^{\mathrm{ID}}_{t}$ and $\bm{u}_{t}=\{P^{\mathrm{DA}}_{t},P^{\mathrm{ID}}_{t}\}$.
  3. (iii)
    First stage: $\Phi_{t}=\Phi_{t}^{(\mathrm{i})}$ and $\bm{u}_{t}=P^{\mathrm{DA}}_{t}$. Second stage: $\Phi_{t}=\Phi_{t}^{(\mathrm{ii})}$ and $\bm{u}_{t}=P^{\mathrm{ID}}_{t}$ for $P^{\mathrm{DA}}_{t}$ fixed.
We assume that $P^{\mathrm{DA}}_{t}$ is allowed to vary sub-hourly. When considering multiple price data clusters ($k>1$), we solve Problem (15) for each cluster $s$, $s\in[1,k]$, using the cluster representative prices $DA_{\lceil t/4\rceil}^{s}$ and $ID_{t}^{s}$. Then, we weight the cost terms $\Phi_{t}^{s}$ based on cluster size by using the respective cluster weights $\lambda_{s}$:
$\mathrm{WDC}=\sum_{s=1}^{k}\frac{24\lambda_{s}}{T}\sum_{t=1}^{T}\Phi_{t}^{s}\,$, (16)
This weighted daily cost (WDC) is then used to compare the different setups.
6.2 Price scenario generation
We investigate different price scenarios, namely the proposed Nominal and Unscaled profiles, and scenarios from $k$-means clustering, $k$-medoids clustering, and hierarchical clustering. Following Kotzur et al. (2018) and Xu et al. (2016), the DA clustering criteria are either (a) the daily mean of the DA prices or (b) the equally-weighted daily mean and daily standard deviation of the DA prices. For the hierarchical clustering, we consider both medoids and centroids as cluster representators. The corresponding generation methodologies are termed “hierarchical-m” and “hierarchical-c”, respectively. In market setups (ii) and (iii), we cluster regarding (c) the daily mean of the DA prices and the daily standard deviation of the ID-DA market deviation, again equally weighted. This clustering strategy is based on previous findings on price component correlation and frequency analysis (Rahimiyan and Baringo, 2016; Germscheid et al., 2022). Here, we focus on comparing the DR case study results for different price generation methods rather than a detailed analysis of the clustering techniques. Therefore, a limited number of clustering configurations is considered, omitting optimization over a different number of clustered profiles (or clusters) within the same clustering technique. Further exploration of clustering configurations for price generation in a DR setup is of interest and left for future work. In total, we here analyze 63 price profiles in 33 price scenarios, numbered S1 to S33, as detailed in Tables 6 to 8.
6.3 Implementation
We use the Python implementation of the $k$-means, $k$-medoids and hybrid clustering algorithms available in the sklearn (Pedregosa et al., 2011) package. The number of clusters considered results from application of the elbow method implemented in the python package kneed (Satopaa et al., 2011). We implement Problem (15) in GAMS and solve the problem using CPLEX with default solver settings. Similar to Section 5, we use historic DA and ID3 price data of the year 2023 from EPEX Spot.
6.4 Results
We discuss the case study results for the market setups (i) to (iii). Based on these results, we extract general conclusions in Section 6.5.
6.4.1 Setup i: Scheduling on DA market
Table 6 (upper part) shows that all scenario generation techniques produce results close to the WDC baseline of 5128 €. The maximum (absolute) deviation of $2.1\,\mathrm{\%}$ occurs in S8, meaning that the schedule underestimates the average energy costs of one year of operation. In general, multi-criterion clustering outperforms single-criterion clustering, which highlights the importance of considering both standard deviation and absolute prices. The schedules obtained when using price scenarios from $k$-means clustering (S4), hierarchical clustering with centroids (S10 and S11), and the Unscaled profile (S3) result in higher costs than for the full year (S1). We explain this loss by the data smoothing associated with profile averaging. The results confirm literature observations that $k$-medoids outperform other clustering techniques (Kotzur et al., 2018; Teichgraeber and Brandt, 2019). Our Nominal DA profile yields $0.6\,\mathrm{\%}$ WDC error, which is comparable to $k$-medoids.
Fig. 10: Single-day Nominal profile (S2) versus profiles from $k$-medoids clustering (S7).
For visual comparison, we illustrate the single-day profiles of $k$-medoids multi-criterion clustering (S7) and the Nominal profile (S2) in Fig. 10. The Nominal profile lies close to the dominant cluster ($s=4$). The corresponding scheduling power profile results are given in the SI.
In some sense, our average-scale-shift strategy is loosely related to clustering techniques. Hence, to empirically examine similarities and differences, we next compare to single-cluster profiles (lower part of Table 6). Comparing the centroid-based scenarios (S12, S13, S18, S19) and our (non-scaled and non-shifted) Unscaled profile reveals an exact match, which is not surprising due to the averaging of all data for derivation of the single cluster representator. Conversely, single-day scheduling based on our Nominal profile notably outperforms all other results, highlighting its high potential despite simplicity of the method. Additionally, we notice that for $k=1$, the $k$-medoids clustering (S14 and S15) performs significantly worse than $k$-means or hierarchical clustering. Finally, we observe no systematic bias in WDC deviation between single- and multiple-profile generation. The absence of bias stems from variation in clustering parameters leading to profiles that may or may not capture outliers of the full-year price data. We highlight that the existence of price outliers, combined with active path constraints and periodical terminal constraints, as outlined in Section 6.1, introduces variability in the results. In a more complex DR setup or a nonconvex DR optimization problem, this variability becomes even more pronounced, leading to less predictable deviation in the results.
Table 6: Results of Case i. The percentages state the relative deviation of WDC from the benchmark (S1). Clustering criteria (a) and (b) as described in Section 6.2.
Generation Crit. k WDC (€) Dev. (%)
S1 Full year - - 5128 $\phantom{-\,}0.0$
S2 Nominal - - 5099  0.6-\,0.6
S3 Unscaled - - 5193 $\phantom{-\,}1.3$
S4 $k$-means (a) 4 5186 $\phantom{-\,}1.1$
S5 $k$-means (b) 4 5186 $\phantom{-\,}1.1$
S6 $k$-medoids (a) 5 5180 $\phantom{-\,}1.0$
S7 $k$-medoids (b) 4 5129 $\phantom{-\,}2\cdot 10^{-2}$
S8 Hierarchical-m (a) 4 5021  2.1-\,2.1
S9 Hierarchical-m (b) 4 5067  1.2-\,1.2
S10 Hierarchical-c (a) 4 5185 $\phantom{-\,}1.1$
S11 Hierarchical-c (b) 4 5188 $\phantom{-\,}1.2$
S12 $k$-means (a) 1 5193 $\phantom{-\,}1.3$
S13 $k$-means (b) 1 5193 $\phantom{-\,}1.3$
S14 $k$-medoids (a) 1 5299 $\phantom{-\,}3.3$
S15 $k$-medoids (b) 1 5423 $\phantom{-\,}5.7$
S16 Hierarchical-m (a) 1 5248 $\phantom{-\,}2.3$
S17 Hierarchical-m (b) 1 5110  0.4-\,0.4
S18 Hierarchical-c (a) 1 5193 $\phantom{-\,}1.3$
S19 Hierarchical-c (b) 1 5193 $\phantom{-\,}1.3$
6.4.2 Setup ii: Simultaneous scheduling on DA & ID market
We present the scheduling results in Table 7. As expected, simultaneous participation in the DA and ID markets reduces costs compared to the DA-only setting. The cost deviations from the baseline are higher (up to 9.4 %) than in Case (i), with the largest errors for centroid-based clustering (S23, S26) and our Unscaled profile (S22). We attribute the higher deviations to a high error sensitivity in the interplay of two price profiles. Again, we observe substantial deviations for centroid-based clustering (S23, S26) and our Unscaled profile (S22). In contrast, using our Nominal DA and ID profiles (S21) only yields slightly larger errors than the medoid-based ($k=6,4$) multi-scenarios (S24 and S25), and give an absolute error of 0.6 % over scheduling based on the full year profiles.
Table 7: Results of Case ii. The percentages state the relative deviation of WDC from the benchmark (S20). S23 to S26 use clustering criteria (c) as described in Section 6.2.
Generation k WDC (€) Dev. (%)
S20 Full year - 4393 $\phantom{-\,}0.0$
S21 Nominal - 4318 1.7-1.7
S22 Unscaled - 4798 $\phantom{-\,}9.2$
S23 $k$-means 5 4805 $\phantom{-\,}9.4$
S24 $k$-medoids 6 4367 0.6-0.6
S25 Hierarchical-m 4 4369 $\phantom{-\,}0.6$
S26 Hierarchical-c 4 4808 $\phantom{-\,}9.4$
6.4.3 Setup iii: Two-stage scheduling on DA & ID market
Table 8 collects the scheduling results of the two-stage problem. Here, the values of WDC and relative deviation fall between Cases i and ii. The best results are obtained when using our Nominal DA and ID profiles (S28), closely followed by the medoid-based hierarchical method (S33) and $k$-medoids clustering (S31). As before, centeroid-based clustering (S30 and S33) and our Unscaled profiles (S29) provide the least accurate prediction of energy costs. We found similar results when considering a one-week horizon, i.e., $T=672$ (see SI).
Table 8: Results of Case iii. The percentages state the relative deviation of WDC from the benchmark (S20). S30 to S33 apply clustering criteria (c) as described in Section 6.2
Generation k WDC (€) Dev. (%)
S27 Full year - 5029 $\phantom{-\,}0.0$
S28 Nominal - 4996 0.6-0.6
S29 Unscaled - 5152 $\phantom{-\,}2.4$
S30 $k$-means 5 5136 $\phantom{-\,}2.1$
S31 $k$-medoids 6 4968 1.2-1.2
S32 Hierarchical-m 4 5064 $\phantom{-\,}0.7$
S33 Hierarchical-c 4 5135 $\phantom{-\,}2.1$
6.5 Discussion
Throughout all scheduling setups (i) to (iii), we have calculated cost errors for the consideration of the proposed price generation and clustering techniques to the use of full-year data. We have shown that employing a single Nominal profile scenario facilitates accurate prediction and generalization of economic savings within a DSM scheduling strategy. Notably, in Case i, using our Nominal scenario outperformed state-of-the-art $k=1$ clustering methods (except hierarchical-m clustering with similar results). More generally, the predictions with our price scenario generation strategy consistently belongs to the two most accurate result throughout all cases, having generalization errors below ± $\pm\,2\,\mathrm{\%}$. This finding supports the use of (variance) scaling within profile construction.
Considering that DSM scheduling savings commonly lie between 2 % and 10 % (Otashu and Baldea, 2018; Caspari et al., 2020; Mucci et al., 2023), we regard scenario generalization errors above 2 % as economically prohibitive. Hence, our scenario generation method produces acceptable generalization errors in the present study. This indicates that a single representative profile can be sufficiently descriptive for early-stage assessment of processes and operation strategies even outperforming multi-scenario clustering methods.
Among all clustering techniques, the medoid-based methods (namely $k$-medoids and hierarchical clustering with a medoid-derived center profile) produce the most accurate scheduling results regarding generalization from a single day to the entire year. Further, considering multiple (here two) clustering criteria over a single criterion yields superior results. Both observations agree with the literature (Teichgraeber and Brandt, 2019; Kotzur et al., 2018).
When considering a multi-stage problem (iii), the Nominal profile proves to be more effective than clustering techniques. In that case, multi-criteria clustering struggles with complex criteria selection and inconsistent representation of individual market behaviors. On the other hand, clustering for each market separately would disregard crucial dependencies between the prices as discussed by Schäfer et al. (2020), possibly preventing conclusions about the proposed DR method. Our method overcomes these challenges by retaining key statistical values, namely the average of the DA profile, the standard deviation of both DA and ID profiles, and the dominant fluctuation patterns of both profiles. Moreover, considering the DA-ID deviation allows for more substantial conclusions and facilitates a zero-integral constraint on the constructed price tuple.
@@ 7  Conclusions
We present a method to construct electricity price scenarios based on an averaging and scaling procedure using historical price data. Our approach is capable of generating both DA and ID scenarios. To reassure consistency, we adjust the cumulative deviation of a DA-ID pair to zero-mean deviation. Our method is constructed to be simple and intuitive, to have few adjustable parameters, and thus be easily reproducible. At the same time, the method can provide sufficiently realistic price profiles, by utilizing recent price data. In particular, we showed that the Nominal profile generation strategy facilitates the generalization from a single-day scheduling horizon to longer time periods, e.g., a full year. Overall, we aim to ease the selection of a price profile in future case studies on DSM.
We have applied our method to historical DA and ID data from EPEX spot for the year 2023. Therein, we generated multiple price scenarios with varying magnitude of price fluctuation. These profiles are ready-to-use and openly accessible [^3] https://git.rwth-aachen.de/avt-svt/public/representative-electricity-price-profiles. The suggested DA and ID profiles closely represent historical time-series of DA and ID profiles realized on the same day. In a DR case study considering DA and ID markets, we compared the suggested price scenarios to single- and multi-clustering techniques. Therein, our suggested Nominal price scenarios provided consistently accurate prediction, comparable to medoid-based multi-scenario hierarchical clustering, while outperforming all other clustering methods. Further, we have recently presented the first application of the price scenarios in a more complex integrated scheduling and control case study Schulze et al. (2024).
We acknowledge that some advanced scenario generation methods additionally consider higher-order statistical moments in their generation procedure and might lead to even more accurate results. However, these approaches are inherently more involved, result in a greater number of adjustable parameters, and can trigger considerably high computational times. In general, both historical data-based and scenario forecasting methods have been inaccurate at predicting the long-term evolution of prices. Clearly, this long-term evolution is subject to many factors, including trading policy and political changes. The forecast prices additionally refer to a specific forecasting time period and do not serve for generalizable results, e.g., they are time-specific and do not provide an average basis for evaluating the broader applicability when benchmarking a new DR concept.
@@
Hence, we focus on approaches to reflect the current character of the electricity price on spot markets. Since we are interested in an intuitive and deterministic method rather than covering all possible price profile shapes, we accept to trade such a flexibility for the desired simplicity of the proposed averaging-based method.
@@
Acknowledgements. The authors gratefully acknowledge the financial support of the Kopernikus project SynErgie by the Federal Ministry of Education and Research (BMBF), and the project supervision by the project management organization Projektträger Jülich (PtJ). The authors gratefully acknowledge the financial support of and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)—333849990/GRK2379 (IRTG Hierarchical and Hybrid Approaches in Modern Inverse Problems). Moreover, we thank Simone Mucci for fruitful discussions as well as Sonja Germscheid and Eike Cramer for constructive feedback on the manuscript.